Hello, I'm completely stuck on this question and have tried to find help in my textbook and notes but are unsuccessful. The question is
6. A bathtub which already had 25L in it is filling with warm water at the rate of 5L/minutes.
A- construct a rule for the bathtubs volume, V (L), in terms of time, t (mins).
My answer is V=25+5t
B- how much water will be in the bathtub after 7 minutes?
My answer is V=25+5(7), V=25+35, V=60
C- if the bathtub can only hold 250L, how long will it take before it is dull?
So I'm stuck on part C, I think I'd have to put the 250l= as the start of the equation but after this I'm lost, would you be able to help me please, thank you 🙂

Answer by Alan3354(69443)

6. A bathtub which already had 25L in it is filling with warm water at the rate of 5L/minutes.
A- construct a rule for the bathtubs volume, V (L), in terms of time, t (mins).
My answer is V=25+5t
B- how much water will be in the bathtub after 7 minutes?
My answer is V=25+5(7), V=25+35, V=60
---------------
C- if the bathtub can only hold 250L, how long will it take before it is dull?
------
V = 25 + 5t
250 = 25 + 5t
5t = 225
t = 45 minutes
=====================
A & B are correct